Balance Sheet (Millions of S) Assets Cash and securities Accounts receivable Inventories Total current assets Net plant and e what is the debt to campital ratio?

and ROIC

and dividends per share
Answer #1
Debt to capital ratio = (Debt/Debt + Shareholders equity)
Debt = Total liabilities
Shareholders equity = Total Assets - Total Liabilities
Total Liabilities = 53900
Total Assets = 70000
Debt to capital ratio = (53900)/((53900+(70000 - 53900))
Debt to capital ratio = (53900)/(53900+16100)
Debt to capital ratio = (53900/70000)
Debt to capital ratio = .77.
The debt to capital ratio is equal to 77%.
Return on invested capital (ROIC) = (Net income - Dividends)/Total capital
Total capital = Total Assets = Total Liabilities + Shareholders equity
$ millions
Net income 2538
Dividends 888.3
Total capital 70000
Total shares outstanding 500 million
ROIC = (2538 - 888.3)/70000
ROIC = .023567
The ROIC is 2.36%.
The dividend per share = Sum of dividends over a period/Ordinary shares outstanding for the period
The dividend per share = 888.3/500
Dividend per share = 1.7766
The dividend per share is $1.78.